Ask your own question, for FREE!
Mathematics 11 Online
OpenStudy (anonymous):

The forecast calls for 30% chance of snow today and a 40% chance of snow tomorrow. What are the chances it will snow on these 2 days in a row?

OpenStudy (anonymous):

These are independent events, so you just multiply probabilities according to the multiplication rule.

OpenStudy (anonymous):

@goxenul So i should multiply both numbers by 2?

OpenStudy (anonymous):

noo....0.3*0.4

OpenStudy (ganpat):

no.. :) what he is telling is multiply 0.3 * 0.4 = 0.12... that is the probability..

OpenStudy (anonymous):

Ohhh thank you guys!

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!